### v1

@@ Methods
The typical XRD images for the S2 and S4 single-crystal samples are in Fig. S1c) and d) in the supplemental material. The XRD patterns for powder samples are in Fig. S1a) and b) in the supplemental material.
First, we index the powder XRD peaks with hkl () and the calculated lattice constant c are 25.4 Å for the S2 and 58.2 Å for the S4 samples, respectively. Those values are much larger than the typical values 6.7 Å of pure graphite, suggesting intercalation by -RuCl3. We estimated the number of RuCl3 layers and figured out the stacking sequence of the graphene and -RuCl3 layers in a unit cell based on the analogy with FeCl3, CrCl3, and AlCl3, [Dresselhaus2002, Cowley1956, Stumpp1977] We adopted the following assumptions.
@@
Fig. S1a) and b) in the supplemental material show the hkl indexing for the single crystal samples. Note that all the observed peaks are for hkl (l0) due to the large c-axis and limited sample oscillation angle during the XRD measurements. Based on the indexing results for powder and single crystal samples and the assumptions mentioned above, we conclude the crystal structure for the sample as shown in the Table 1and Fig. 1b), c) and d).
@@
We perform density functional theory (DFT) calculations as implemented in VASP [Kresse_VASP] simulation package version 6.3.0. In all calculations we apply the Generalized Gradient Approximation (GGA) [GGA] as exchange-correlation functional. We use the following pseudopotentials in the projector augmented wave method (PAW) [blochl_projector_1994, kresse_ultrasoft_1999] Ru_pv, Cl and C provided by the VASP package. In all calculations the basis set plane-wave cut-off for the expansion is set to 500 eV. In all crystal structures we relax the -RuCl3 layer and keep the graphene layers geometry fixed. The graphene--RuCl3 layer distance is kept fixed at approximately the average experimental value of Å, the adjacent graphene sheets at Å, and the C-C distance in each graphene layer is Å, consistent with previous reports [Chen2011_tight_binding_graphene]. The relaxations are performed via the -point version of VASP with a convergence criterion of the forces for each atom in the -RuCl3 layer in each direction of 0.005 eV/Å assuming a non-magnetic state. For the crystal cell optimization we include van der Waals corrections via the DFT-D3 method of Grimme [grimme2010consistent, grimme2011effect].
Further electronic properties are calculated on a 662 k-mesh. The supercell electronic band structures are unfolded into the primitive Brillouin zone (BZ) by using the KPROJ software [chen2025kproj]. The Fermi surfaces are generated from sequences of unfolded electron band structures. The charge transfer was estimated via the Bader analysis [HENKELMAN2006354], and k-mesh convergence has been checked. The spatial resolved charge transfer was calculated via the VASPKIT package [VASPKIT]. The quantum oscillation frequencies are calculated from the unfolded graphene layers Fermi surfaces. Due to the smeared Fermi surfaces as a consequence of the unfolding, the Fermi pocket boundaries are taken from the highest intensity points and connected to form a closed area. Therefore, all reported frequencies inherit a small systematic error which does not change the conclusions of the study.
Additionally, as a cross-check we compute the quantum frequencies for the AA structure via the C++ program dhva [backes2014electronic, numpy2024], which implements the dHvA frequency extraction algorithm introduced by Rourke and Julian [julian2012numerical]. Calculations are performed on a supercell grid comprising k-points distributed over 64 unit cells. Higher k-point density is obtained through tricubic interpolation [lekien2005tricubic]. The magnetic field is applied in the (001) direction.
@@ Acknowledgements
Data availability statement
The datasets generated during the current study are avail- able upon reasonable request.
@@
See pages 1 of Supplement_V1_arxiv.pdfSee pages 2 of Supplement_V1_arxiv.pdfSee pages 3 of Supplement_V1_arxiv.pdfSee pages 4 of Supplement_V1_arxiv.pdf

### v2

@@ Methods
The typical XRD images for the S2 and S4 single-crystal samples are in Figures S1c and d in the supplemental material. The XRD patterns for powder samples are in Figures S1a and b in the supplemental material.
First, we index the powder XRD peaks with hkl () and the calculated lattice constant c are 25.4 Å for the S2 and 58.2 Å for the S4 samples, respectively. Those values are much larger than the typical values 6.7 Å of pure graphite, suggesting intercalation by -RuCl3. We estimated the number of RuCl3 layers and figured out the stacking sequence of the graphene and -RuCl3 layers in a unit cell based on the analogy with FeCl3, CrCl3, and AlCl3, [28, 29, 30] We adopted the following assumptions.
@@
Figures S1a and b in the supplemental material show the hkl indexing for the single crystal samples. Note that all the observed peaks are for hkl (l0) due to the large c-axis and limited sample oscillation angle during the XRD measurements. Based on the indexing results for powder and single crystal samples and the assumptions mentioned above, we conclude the crystal structure for the sample as shown in the Table 1and Figure 1b, c and d.
@@
We perform density functional theory (DFT) calculations as implemented in VASP [43] simulation package version 6.3.0. In all calculations we apply the Generalized Gradient Approximation (GGA) [44] as exchange-correlation functional. We use the following pseudopotentials in the projector augmented wave method (PAW) [45, 46] Ru_pv, Cl and C provided by the VASP package. In all calculations the basis set plane-wave cut-off for the expansion is set to 500 eV. In all crystal structures we relax the -RuCl3 layer and keep the graphene layers geometry fixed. The graphene--RuCl3 layer distance is kept fixed at approximately the average experimental value of Å, the adjacent graphene sheets at Å, and the C-C distance in each graphene layer is Å, consistent with previous reports [47]. The relaxations are performed via the -point version of VASP with a convergence criterion of the forces for each atom in the -RuCl3 layer in each direction of 0.005 eV/Å assuming a non-magnetic state. For the crystal cell optimization we include van der Waals corrections via the DFT-D3 method of Grimme [48, 49].
Further electronic properties are calculated on a 662 k-mesh. The supercell electronic band structures are unfolded into the primitive Brillouin zone (BZ) by using the KPROJ software [50]. The Fermi surfaces are generated from sequences of unfolded electron band structures. The charge transfer was estimated via the Bader analysis [51], and k-mesh convergence has been checked. The spatial resolved charge transfer was calculated via the VASPKIT package [52]. The quantum oscillation frequencies are calculated from the unfolded graphene layers Fermi surfaces. Due to the smeared Fermi surfaces as a consequence of the unfolding, the Fermi pocket boundaries are taken from the highest intensity points and connected to form a closed area. Therefore, all reported frequencies inherit a small systematic error which does not change the conclusions of the study.
Additionally, as a cross-check we compute the quantum frequencies for the AA structure via the C++ program dhva [53, 54], which implements the dHvA frequency extraction algorithm introduced by Rourke and Julian [55]. Calculations are performed on a supercell grid comprising k-points distributed over 64 unit cells. Higher k-point density is obtained through tricubic interpolation [56]. The magnetic field is applied in the (001) direction.
@@ Acknowledgements
Author Contributions
N.H. and M.C synthesized the crystals. S.M. performed the transport and quantum oscillation measurements. T.M. and M.C. performed the X-ray experiments. A.R. performed the ab initio simulations. R.V. and D.M. conceived and supervised the project. All authors contributed to the discussions and writing of the manuscript.
@@
Data availability statement
The datasets generated during the current study are avail- able upon reasonable request. not applicable)) The authors declare no conflict of interest.